# CodeTool: Enhancing Programmatic Tool Invocation of LLMs via Process Supervision

### v1

@@ Abstract
Tool invocation significantly enhances the capabilities of Large Language Models (LLMs), yet challenges persist, particularly in complex task scenarios. Current methods, such as instruction-enhanced reasoning and supervised fine-tuning, often result in unnecessarily long reasoning paths and face difficulties in verifying the correctness of intermediate steps. In this paper, we propose CodeTool, a novel framework for stepwise code generation that improves LLM tool invocation by leveraging the concise and easily verifiable nature of code. CodeTool incorporates two distinct process rewards: the On-the-spot Reward, which provides immediate feedback on the accuracy of each tool invocation, and the Latent Reward, which assesses the contribution of each step toward overall task completion. By maximizing the cumulative reward of the On-the-spot and Latend Rewards at each step, LLMs are guided to follow efficient and accurate reasoning paths. Extensive experiments on StableToolBench and RestBench-TMDB demonstrate the superiority of CodeTool over existing approaches.
11footnotetext: Work done during an internship at Tencent Hunyuan.22footnotetext: Equal Contribution.33footnotetext: Corresponding Authors.
Figure 1: Comparison of tool invocation in JSON format and our proposed stepwise code generation framework: (a) JSON-based invocation is token-heavy and prone to truncation, risking loss of key information. (b) The stepwise code generation framework uses loops to handle request-intensive tasks efficiently, with stepwise supervision ensuring correctness of intermediate steps.
@@ 3.3 Process Reward Supervision
where $\Delta_{correct}$ and $\Delta_{total}$ denote the number of paths that are correctly executed and the total number of executed paths, respectively, from the $t$-th step to its leaf nodes, while $s_{1:t}$ denotes the reasoning sequence from the first step to step $t$.
@@ 3.4 Process Latent Reward Model Training
For model training, we select user queries from the ToolBench (Qin et al., 2023) training set that remain solvable, meaning that the tools or APIs required to solve these queries are still callable, and use them to construct the intermediate process data. To balance computational resource usage and inference time, we employ a depth-first search algorithm to construct an action search space resembling a binary tree for each user query. Specifically, during the code generation for each tool invocation step, we perform sampling twice. Subsequently, at each hierarchical level of the action search space, we collect process data with varying Latent Reward values, which are regarded as a key indication for comparing the levels of potential. Ultimately, we leverage the data to train the PRM. The entire process is fully automated, ensuring efficiency and requiring no human intervention.
@@ 4.3 Baselines and Evaluation Metric
We conduct experiments using both open-source and closed-source LLMs. For open-source LLMs, we primarily compare our method with the well-established baseline, ToolLLaMA (Qin et al., 2023), which is fine-tuned from LLaMA on successful tool execution chains. Additionally, we include two reinforcement learning-based baselines, TP-LLaMA (Chen et al., 2024) and StepTool (Yu et al., 2024), which utilize direct preference optimization (Rafailov et al., 2024) and proximal policy optimization (Schulman et al., 2017), respectively. To take full advantage of the code generation capability of LLMs and ensure a fair comparison at the same time, we employ CodeLlama-7B [^1] https://huggingface.co/codellama/CodeLlama-7b-hf, Qwen2.5-7B-Instruct [^2] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct and Qwen2.5-Coder-7B-Instruct [^3] https://huggingface.co/Qwen/Qwen2.5-Coder-7B-Instruct as the code generation models for CodeTool. All models share the same number of parameters (i.e., 7B) as the baselines. For closed-source LLMs, we adopt GPT-3.5-Turbo-0613 and GPT-4-Turbo-Preview, each representing different performance levels and capabilities within the GPT series. Following Qin et al. (2023); Yu et al. (2024), we employ two inference strategies for baselines: Chain of Thought (CoT) (Wei et al., 2022) and Depth-First Search Decision Tree (DFSDT) (Qin et al., 2023).
Following StableToolBench, we utilize the Solvable Pass Rate (SoPR) as the evaluation metric. Specifically, GPT-4 (gpt-4-turbo-2024-04-09) is leveraged as the evaluator to categorize the answers into ‘‘Solved’’, ‘‘Unsolved’’, or ‘‘Unsure’’, with corresponding scores of 1, 0.5, and 0, respectively, contributing to the overall SoPR calculation. However, evaluation experiments on StableToolBench have shown that different models exhibit varying preferences regarding the degree to which the final answer resolves the query, leading to unstable evaluation results. To address this, we test the evaluation script provided by StableToolBench and introduce clearer criteria for assessing these three categories (detailed in Appendix B), significantly enhancing the stability of model evaluations.
@@ 4.4 Training Settings
We only need to train the process Latent Reward model. In our experiments, we train such a model using Qwen2.5-7B-Instruct. To avoid disrupting the native structure of the LLM, we adopt a generative PRM training method. Specifically, we designate two special tokens to represent the ‘‘more potential’’ and ‘‘less potential’’ labels based on Latent Reward values, and then fully reuse the training method of SFT. We train the Qwen2.5-7B-Instruct model on the collected process data for 2 epochs with a learning rate of 1e-6.
@@
Table 2: Performance comparison in terms of SoPR between CodeTool and baselines. The best performances are highlighted in bold, while suboptimal ones are marked with underline. We reproduce StepTool on ToolLLaMA using the released code. All results are assessed on the filtered StableToolBench test set using the improved SoPR evaluation prompt (refer to Section 4.3 for details about the performance evaluation prompt).
@@ 5.1 Main Results
Table 2 presents the performance comparison of CodeTool with baselines on both open-source and closed-source LLMs. Below are some key observations drawn from the results: (1) While CodeLlama-7B exhibits lower performance compared to ToolLLaMA when the DFSDT strategy is adopted, the more capable Qwen2.5-7B-Instruct and Qwen2.5-Coder-7B-Instruct, both of which possess stronger coding and instruction-following abilities, achieve significantly higher SoPR scores on the test set, demonstrating the effectiveness of the CodeTool framework. Although Qwen2.5-7B-Instruct and Qwen2.5-Coder-7B-Instruct perform comparably across different test subsets, with each showing advantages in certain areas, Qwen2.5-Coder-7B-Instruct ultimately offers more consistent and overall better performance. (2) For closed-source LLMs, both GPT-3.5-Turbo-0613 and GPT-4-Turbo-Preview demonstrate comparable performance on the SoPR metric when employing the DFSDT strategy with JSON-format-based tool invocations. Notably, GPT-4-Turbo-Preview, with its superior code generation capabilities, further elevates the SoPR metric when paired with CodeTool, enhancing the performance of Qwen2.5-Coder-7B-Instruct by 1.86%. (3) The results on both open-source and closed-source LLMs suggest that the more advanced the model’s coding capabilities, the better its tool invocation performance in terms of the SoPR metric when paired with our CodeTool.
@@ 5.2 Generalization Performance on RestBench-TMDB
ReAct (Yao et al., 2023) 61.00
RestGPT (Song et al., 2023) 65.00
CodeAct (Wang et al., 2024c) 63.00
ToolLLaMA (Qin et al., 2023) 72.00
ATC (Shi et al., 2024) 89.00
CodeTool 92.00
Table 3: Performance comparison in terms of success rate on RestBench-TMDB.
To evaluate the generalization capability of the CodeTool framework, we conduct additional experiments on another widely recognized benchmark, RestBench-TMDB Song et al. (2023), which contains 100 tasks involving 54 tools designed for movie-related scenarios. Notably, we do not retrain the process Latent Reward model for this benchmark; instead, we utilize the model previously trained for StableToolBench directly. We adopt the Success Rate metric provided by RestBench-TMDB for evaluation, which relies on human assessment to ascertain whether the model’s output successfully fulfills the user query. The results, as shown in Table 3, indicate that CodeTool also achieves the best performance, with a Success Rate of 92%, surpassing ATC by 3.37%, which involves writing complete Python code to solve user queries.
@@ Appendix C StableToolBench Test Set Filtering Rules
Appendix D Case Study

### v2

@@ Abstract
Tool invocation significantly enhances the capabilities of Large Language Models (LLMs), yet challenges persist, particularly in complex task scenarios. Current methods, such as instruction-enhanced reasoning and supervised fine-tuning, often result in unnecessarily long reasoning depths and face difficulties in verifying the correctness of intermediate steps. In this paper, we propose CodeTool, a novel framework for stepwise code generation that improves LLM tool invocation by leveraging the concise and easily verifiable nature of code. CodeTool incorporates two distinct process rewards: the On-the-spot Reward, which provides immediate feedback on the accuracy of each tool invocation, and the Latent Reward, which assesses the contribution of each step toward overall task completion. By maximizing the cumulative reward of the On-the-spot and Latend Rewards at each step, LLMs are guided to follow efficient and accurate reasoning paths. Extensive experiments on StableToolBench and RestBench-TMDB demonstrate the superiority of CodeTool over existing approaches. Our implementation is available at LimOkii/CodeTool.
11footnotetext: Work done during an internship at Tencent Hunyuan.footnotetext: Equal Contribution.footnotetext: Corresponding Author.
Figure 1: Comparison of tool invocation in JSON format and our proposed stepwise code generation framework: (a) JSON-based invocation is token-heavy when handling request-intensive tasks, and is prone to truncation, risking the loss of key information. (b) The stepwise code generation framework uses loops to handle request-intensive tasks efficiently, with stepwise supervision ensuring correctness of intermediate steps.
@@ 3.3 Process Reward Supervision
where $\Delta_{correct}$ and $\Delta_{total}$ denote the number of paths that are correctly executed with final answer labeled as ‘‘Solved’’ and the total number of executed paths, respectively, from the $t$-th step to its leaf nodes, while $s_{1:t}$ denotes the reasoning sequence from the first step to step $t$.
@@ 3.4 Process Latent Reward Model Training
For model training, we select user queries from the ToolBench (Qin et al., 2023) training set that remain solvable, meaning that the tools or APIs required to solve these queries are still callable, and use them to construct the intermediate process data. To balance computational resource usage and inference time, we employ a depth-first search algorithm to construct an action search space resembling a binary tree for each user query. Specifically, during the code generation for each tool invocation step, we perform sampling twice. Subsequently, at each hierarchical level of the action search space, we collect process data with varying Latent Reward values, which are regarded as a key indication for comparing the levels of potential. Ultimately, we leverage the data to train the PRM. The entire process is fully automated, ensuring efficiency and requiring no human intervention. Further details can be found in the Appendix E
@@ 4.3 Baselines and Evaluation Metric
We conduct experiments using both open-source and closed-source LLMs. For open-source LLMs, we primarily compare our method with the well-established baseline, ToolLLaMA (Qin et al., 2023), which is fine-tuned from LLaMA on successful tool execution chains. Additionally, we include two reinforcement learning-based baselines, TP-LLaMA (Chen et al., 2024) and StepTool (Yu et al., 2024), which utilize direct preference optimization (Rafailov et al., 2024) and proximal policy optimization (Schulman et al., 2017), respectively. To take full advantage of the code generation capability of LLMs and ensure a fair comparison at the same time, we employ CodeLlama-7B [^1] https://huggingface.co/codellama/CodeLlama-7b-hf, Qwen2.5-7B-Instruct [^2] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct and Qwen2.5-Coder-7B-Instruct [^3] https://huggingface.co/Qwen/Qwen2.5-Coder-7B-Instruct as the code generation models without any additional fine-tuning for CodeTool. All models share the same number of parameters (i.e., 7B) as the baselines. For closed-source LLMs, we adopt GPT-3.5-Turbo-0613 and GPT-4-Turbo-Preview, each representing different performance levels and capabilities within the GPT series. Following Qin et al. (2023); Yu et al. (2024), we employ two inference strategies for baselines: Chain of Thought (CoT) (Wei et al., 2022) and Depth-First Search Decision Tree (DFSDT) (Qin et al., 2023).
Following StableToolBench, we utilize the Solvable Pass Rate (SoPR) as the evaluation metric. Specifically, GPT-4 (gpt-4-turbo-2024-04-09) is leveraged as the evaluator to categorize the answers into ‘‘Solved’’, ‘‘Unsure’’, or ‘‘Unsolved’’, with corresponding scores of 1, 0.5, and 0, respectively, contributing to the overall SoPR calculation. However, evaluation experiments on StableToolBench have shown that different models exhibit varying preferences regarding the degree to which the final answer resolves the query, leading to unstable evaluation results. To address this, we test the evaluation script provided by StableToolBench and introduce clearer criteria for assessing these three categories (detailed in Appendix B), significantly enhancing the stability of model evaluations.
@@ 4.4 Training Settings
We only need to train the process Latent Reward model. In our experiments, we train such a model using Qwen2.5-7B-Instruct. To avoid disrupting the native structure of the LLM, we adopt a generative PRM training method. Specifically, we designate two special tokens to represent the ‘‘more potential’’ and ‘‘less potential’’ labels based on Latent Reward values, and then fully reuse the training method of SFT. We train the Qwen2.5-7B-Instruct model on the collected process data for 2 epochs with a learning rate of 1e-6. More training details can be found in the Appendix E
@@
Qwen2.5-Instruct-7B CoT JSON 56.11 58.20 51.26 54.89 61.00 44.12 54.26
DFSDT JSON 61.07 63.11 62.18 54.89 66.00 55.88 60.52
CodeTool Code 63.74 68.44 65.68 55.98 72.50 58.82 64.19
@@
Table 2: Performance comparison in terms of SoPR between CodeTool and baselines. The best performances in open-source LLMs are highlighted in bold, while suboptimal ones are marked with underline. We reproduce StepTool on ToolLLaMA-v2 using the released code. All results are assessed on the filtered StableToolBench test set using the improved SoPR evaluation prompt (refer to Section 4.3 for details about the evaluation prompt).
@@ 5.1 Main Results
Table 2 presents the performance comparison of CodeTool with baselines on both open-source and closed-source LLMs. Below are some key observations drawn from the results: (1) For open-source LLMs, when Qwen2.5-7B-Instruct is used as the code generation model, CodeTool consistently outperforms both the CoT and DFSDT strategies. (2) For open-source Coder LLMs, while CodeLlama-7B exhibits lower performance compared to ToolLLaMA when the DFSDT strategy is adopted, the more capable Qwen2.5-Coder-7B-Instruct, which possesses stronger coding and instruction-following abilities, achieves significantly higher SoPR scores on the test set, demonstrating the effectiveness of the CodeTool framework. Although Qwen2.5-7B-Instruct and Qwen2.5-Coder-7B-Instruct perform comparably across different test subsets, with each showing advantages in certain areas, Qwen2.5-Coder-7B-Instruct ultimately offers more consistent and overall better performance. (3) For closed-source LLMs, both GPT-3.5-Turbo-0613 and GPT-4-Turbo-Preview demonstrate comparable performance on the SoPR metric when employing the DFSDT strategy with JSON-format-based tool invocations. Notably, GPT-4-Turbo-Preview, with its superior code generation capabilities, further elevates the SoPR metric when paired with CodeTool, enhancing the performance of Qwen2.5-Coder-7B-Instruct by 1.86%. (4) The results on both open-source and closed-source LLMs suggest that the more advanced the model’s coding capabilities, the better its tool invocation performance in terms of the SoPR metric when paired with our CodeTool.
@@ 5.2 Generalization Performance on RestBench-TMDB
ReAct (Yao et al., 2023) 61.00 77.13
RestGPT (Song et al., 2023) 65.00 77.49
CodeAct (Wang et al., 2024c) 63.00 80.91
ToolLLaMA (Qin et al., 2023) 72.00 78.29
ATC (Shi et al., 2024) 89.00 84.71
CodeTool 92.00 91.15
Table 3: Performance comparison in terms of success rate and correct path rate on RestBench-TMDB.
To evaluate the generalization capability of the CodeTool framework, we conduct additional experiments on another widely recognized benchmark, RestBench-TMDB Song et al. (2023), which contains 100 tasks involving 54 tools designed for movie-related scenarios. Notably, we do not retrain the process Latent Reward model for this benchmark; instead, we utilize the model previously trained for StableToolBench directly. We adopt the Success Rate and Correct Path Rate metrics provided by RestBench-TMDB for evaluation. Success Rate relies on human assessment to ascertain whether the model’s output successfully fulfills the user query, while Correct Path Rate measures the proportion of ground truth tools in model-generated tool invocations. The results, as shown in Table 3, indicate that CodeTool achieves the best performance, with a Success Rate of 92% and a Pass Rate of 91.15%, surpassing ATC by 3.37% and 7.60%, respectively, which involves writing complete Python code to solve user queries.
@@ 5.3 Analyses
Efficiency Comparison with JSON-based Tool Invocation
To further investigate the differences in reasoning efficiency between code-based and JSON-based methods, particularly in terms of reasoning depth and token consumption, we conduct additional comparative experiments on I2-Category. We use Qwen2.5-7B-Instruct as the code generation model. As shown in Figure 3, in terms of average reasoning depth, the CodeTool framework, compared with JSON-based tool invocation methods that employ CoT and DFSDT strategies, enables both tool invocation and response acquisition within a single interaction step, thereby effectively reducing the number of interactions by nearly half. As shown in Figure 4, the CodeTool framework maintains a comparable level of token consumption to the DFSDT strategy, while outperforming both DFSDT and CoT in terms of SoPR, further underscoring the effectiveness of our approach.
Figure 3: Average reasoning depths of three different strategies across six test subsets in StableToolBench.
Figure 4: Token consumption of three different strategies across six test subsets in StableToolBench.
@@ Appendix C StableToolBench Test Set Filtering Rules
Appendix D Total Number of APIs Involved in the Filtered Test Set
The total number of APIs involved in the filtered test set is shown in Table 8.
I1-I I1-C I1-T I2-I I2-C I3-I
Task 131 122 119 92 100 17
Candidate API 631 271 371 536 341 31
Relevant API 288 171 199 197 172 19
Table 8: The total number of APIs involved in the filtered test set. C, I, T stand for the Category, Instruction, and Tool subgroup of the test set.
Appendix E Details of PRM Training
In the process data collection phase, we draw inspiration from the sampling strategy of MCTS. However, considering the high time and computational cost associated with performing multiple rollouts in MCTS, we adopt a simplified implementation. As described in Section 3.4 of the paper, to balance computational resources and inference efficiency, we employ a Depth-First Search (DFS) algorithm. Specifically, during the code generation phase at each step of tool invocation, we perform two samples for each node’s next action, thereby expanding a binary tree-like search space for code generation paths. For each node in the search tree, we first compute its raw Latent Reward using Equation 4, where a key challenge lies in determining whether the current reasoning path successfully solves the user’s query. In the scenario of code-based tool invocation, we adopt the practice of Shi et al. (2024) by explicitly instructing the model in the prompt to output a complete Final Answer via a print statement at the final step of reasoning. In relatively straightforward tool-invocation scenarios (e.g., numerical queries), the model typically follows this instruction well, accurately providing a complete Final Answer. However, in more complex tasks, it is difficult for the model to directly generate a complete answer in the final step. Thus, we implement a more flexible strategy: after each tool invocation, the model immediately prints out the result returned from that request, as specified explicitly in Figure 6. Subsequently, we pair the model’s thought at each reasoning step with the corresponding tool response, and combine these pairs with the original user query to prompt the LLM to reorganize them into a coherent natural-language text as the Final Answer. After obtaining the Final Answer, we apply the same evaluation prompt used in the SoPR experiment (detailed in Table 6). Based on how well the model-generated Final Answer addresses the user’s original query, LLM classifies each answer into one of three categories: ‘‘Solved’’, ‘‘Unsure’’, or ‘‘Unsolved’’. Accordingly, we automatically assign its final Latent Reward using Equation 5. Leveraging these reward signals, we further collect process data with varying Latent Reward values from every level of the search tree, thereby constructing training samples for the PRM that capture the differences in potential among various paths.
In the training phase of the PRM, to avoid interfering with the structure of the LLM, we follow a generative training strategy similar to Wang et al. (2024b). Specifically, for two child nodes sampled from the same parent node, we assign labels ‘‘more potential’’ and ‘‘less potential’’ based on their Latent Reward values (e.g., using two special tokens ‘‘yes’’ and ‘‘no’’ as supervision signals) for supervised fine-tuning.
During inference, to estimate the Latent Reward of the current step, we compute the logits of the ‘‘yes’’ and ‘‘no’’ tokens output by the model and normalize them using the following formula:
Latent $\text{Latent Reward}=\frac{\text{logit}_{\text{yes}}}{\text{logit}_{\text{yes}}+\text{logit}_{\text{no}}}.$ (8)
Appendix F Analysis of Rewards Conflict
To further investigate the Reward Conflict phenomenon during inference, we conduct inference using Qwen2.5-7B-Instruct as the code generation model on the I2-Category subset, with the number of candidate actions set to 2. In total, we collect 497 pairs of intermediate reasoning data. The detailed reward distribution is shown in Table 9.
Case 1 refers to Both candidates have an On-the-spot Reward of 1 (i.e., both are executable);
Case 2 refers to Both candidates have an On-the-spot Reward of 0 (i.e., neither candidate is executable);
Case 3 refers to One candidate node is executable while the other is not, and the executable node has a higher Latent Reward;
Case 4 refers to One candidate is executable, but has a lower Latent Reward than the non-executable one (i.e., the reward conflict scenario.)
Case Type Number Ratio
Total Samples 497
Case 1 363 73.04%
Case 2 59 11.87%
Case 3 41 8.25%
Case 4 (conflict) 34 6.84%
Table 9: Distribution of case types including reward conflict cases.
We observe that reward conflict (Case 4) is indeed rare, accounting for only 6.84 % of all sampled cases. This low occurrence suggests that, in most situations, the On-the-spot Reward and Latent Reward are aligned.
Appendix G Case Study